Answer the following questions which relate to aggregate expenditure models. a. If C is D100, I is D50, Xn is D-10, G is D30, what is the economy’s equilibrium GDP? b. If real GDP in an economy is currently D200, C is D100, I is D50, Xn is D-10, G is D30. Will the economy’s real GDP rise, fall or stay the same c. Suppose that full employment output in an economy is D200, C is D150, I is D50, Xn is D-10, G is D30. What will be the macroeconomics result? 6


1
Expert's answer
2021-06-07T07:07:02-0400

Answers:

(i) Equilibrium GDP; Y = AE

AE (Aggregate expenditure) = C + I + G + (X-M)

= D(100 + 50 + 30 -10)

= D170


(ii) Determination of GDP deflator = "\frac {Nominal GDP}{Real GDP}\times100" = "\frac{170}{200} \times100" = 85% 0 r 0.85

In the base year, deflator is 100% while current year it is 85%, implies that the Real GDP is likely to fall.


(iii) Economic result:

Real GDP for the current year = D170

Nominal GDP = D(150 + 50 + 30 - 10) = D220

GDP Deflator = "\frac{220}{170} \times100"

= 129.41%

The economic impact is that the real GDP will increase by 29.41%.